# Conditioning diffusion models by explicit forward-backward bridging

### v1

@@ Appendix B Experiment settings
B.1 Common settings
Unless otherwise stated, our Gibbs-CSMC method applies the explicit backward sampling in Section 3.1 and the killing conditional resampling (Karppinen et al., 2023), and our PMCMC method applies the PCN proposal of Cotter et al. (2013) and stratified resampling (Chopin and Papaspiliopoulos, 2020, Chap. 9).
For image-related experiments, we use a standard UNet commonly used in denoising diffusion models (see Ho et al., 2020) with three downsampling and three upsampling layers, and we use the pixel shuffle method for upsampling in the UNet. The initial number of convolution features is 64, followed by 128, and 256 in the three layers. The detailed construction of the neural network is given in our companion code repository at https://github.com/zgbkdlm/fbs.
We randomly split the MNIST dataset into 60,000 training and 10,000 test data points. As for the CelebaHQ dataset (Liu et al., 2015), we split it into 29,000 training and 1,000 test data points. All the image data are normalised within the range $[0,1]$.
All experiments are implemented in JAX (Bradbury et al., 2018) with the Flax (Heek et al., 2023) neural network backend. The experiments are conducted on an NVIDIA A100 80G GPU and the code required to reproduce our empirical findings can be found at https://github.com/zgbkdlm/fbs.
B.2 High-dimensional synthetic conditional sampling
@@
While the CSGM method needs the likelihood distribution $p(y\mid x_{t})$ to work (Song et al., 2021), in this Gaussian special case, the likelihood is analytically tractable, and hence the method can be applied exactly. For the twisted PF, we apply the GP observation model as the twisting function as per Wu et al. (2023, Eq. (8)).
@@
Figure 6 shows the trace of the PMCMC chain of one marginal variable. We see that the parameter $\delta$ affects the autocorrelation of the MCMC chain, which is also reflected in Table 4.1 and Figure 2. The chain appears to be more effective when $\delta$ is small, meaning that the measurement path $t\mapsto Y_{t}$ in Section 3.2 ideally should not move too much. Furthermore, we also see that the chain seems to gradually degenerate as the chain iteration goes for larger moves in the $Y$ space. This is a well-known defect of pseudo-marginal methods (Andrieu and Roberts, 2009) in general, and comes from the fact that proposed paths are likely to be accepted originally due to the variance in the acceptance ratio, and then unlikely once the process has stabilised if the proposed path is too far from the current one. Solving this problem is an interesting avenue of future work and likely involves redesigning the proposal kernel for the path $Y$ to take into account the specific structure of the model.
B.3 Conditional sampling in non-separable noising processes
@@
To show the effect of using a non-separable noising process, we conduct quantitative and qualitative experiments using Schrödinger bridges. In the quantitative experiment, we apply the same GP model in Appendix B.2, but we scale the observation noise covariance to $0.1$ to increase the correlation between $X$ and $Y$. We use a Wishart distribution to randomly generate the reference covariance matrix $Q=q\otimes q$, where $\otimes$ is the outer product, and $q\sim\mathcal{N}(0,I_{d})$. We then consider a Gaussian Schrödinger bridge between $\pi(x,y)$ and $\mathcal{N}(0,Q)$, with reference process $\mathop{}\!\mathrm{d}Z_{t}=\mathop{}\!\mathrm{d}W_{t}$ given as the standard Brownian motion, the formulation of which is analytically tractable (Bunne et al., 2023).
For the qualitative experiments on MNIST images, we train a Schrödinger bridge with reference process $\mathop{}\!\mathrm{d}Z_{t}=-0.5\,\beta_{t}\,Z_{t}\mathop{}\!\mathrm{d}t+\sqrt{\beta_{t}}\mathop{}\!\mathrm{d}W_{t}$ at $t\in[0,0.5]$, where $\beta_{t}\coloneqq(b_{\mathrm{max}}-b_{\mathrm{min}})\,/\,(T-t_{0})\,t+(b_{\mathrm{min}}\,T-b_{\mathrm{max}}\,t_{0})\,/\,(T-t_{0})$, and we set $b_{\mathrm{max}}=5$, $b_{\mathrm{max}}=0.02$, $t_{0}=0$, and $T=0.5$. The reference distribution $\pi_{\mathrm{ref}}$ at $T$ is a standard unit Gaussian. The number of Schrödinger bridge iterations is 20. At each Schrödinger bridge iteration, we run 20 epochs training the forward and backward models, with spatial batch size 64 and temporal batch size 32. We use the Adam (Kingma and Ba, 2015) optimiser with learning rate decaying from $2\times 10^{-4}$ to $2\times 10^{-6}$ in a cosine curve (Loshchilov and Hutter, 2017). For more details on the algorithm used to train the bridge, see De Bortoli et al. (2021, Algorithm 1).
@@
As shown in Figure 7, the initialisation of $X_{0}$ indeed affects the quality of PF samples due to the non-separability of the Schrödinger bridge, while our Gibbs-CSMC sampler returns valid samples. This visible bias stems from the fact that, under non-separability, the choice of $X_{0}$ has a direct impact on the distribution of the path $Y_{1:K}\mid Y_{0}=y,X_{0}$, as conditional independence is then not verified anymore. For the path $Y_{1:K}\mid\{Y_{0}=y,X_{0}\}$ to be marginally distributed according to the forward noising diffusion (as required by Trippe et al., 2023), one would need the initial value of $X_{0}$ to be distributed according to $\pi(x\mid y)$, which is the problem we are trying to solve and is therefore not feasible in practice. While it is reasonable to assume that approximate samples from $\pi(x\mid y)$ can be obtained via other methods, this would incur additional computation and still not eliminate the bias. On the other hand, our Gibbs-CSMC method provides a natural way to produce samples that iteratively converge to $\pi(x\mid y)$, explaining its improved performance.
@@
B.4 Image inpainting and super-resolution
@@
We first pre-train the unconditional model to approximate the resulting score function using the standard denoising score matching method of Song et al. (2020). The optimiser we used for training is Adam with a decaying cosine learning rate schedule, i.e., the learning rate starts from $2\times 10^{-4}$ and ends at $2\times 10^{-6}$ following a cosine curve (Loshchilov and Hutter, 2017). The spatial and temporal batch sizes are 256. We also apply gradient clipping and exponential moving average (with 0.99 decay rate per two iterations) as suggested in Song and Ermon (2020). We run the training for 3,000 epochs and use the model obtained in the last epoch.
In the inpainting task, we apply rectangles of sizes 15 and 32 for MNIST and CelebAHQ, respectively, spawning at random locations (uniformly) in the image. As for super-resolution, we use rates 4 and 2 for MNIST and CelebAHQ. Precisely, rate 2 means that for an image of size 32 by 32, we partition it into 16 by 16 squares, where each square has 2 by 2 pixels; at each square, we randomly select one pixel for the low-resolution image. For each test (inpainting/super-resolution) image, we generate 100 restored image samples, and report the statistics as in Tables 3 and 3. In the table, the LPIPS score uses an Alex net (Zhang et al., 2018). The commonly used Fréchet inception distance is not applied here, since it is not clear how to use it to benchmark MCMC samples (cf. the arguments in Cardoso et al., 2024).
For the CSGM method, we implement the conditional drift as per Song et al. (2021, Sec. I.2). As for the TPF, we also implement the same twisting function as in Wu et al. (2023, Sec. 3.3).
We show samples of inpainting and super-resolution tasks in Figures 912, to qualitatively compare the methods. We find from these figures that our Gibbs-CSMC and PMCMC methods indeed outperform all other methods, in terms of quality and correctness. By correctness, we mean that the samples generated by our methods are realistic and closer to the true images. As an example, in the first panel of Figure 9, the MCMC chains generate a variety of digits “4” which indeed correspond to the true digit, while other methods can generate wrong, albeit more diverse digits.

### v2

@@ Appendix E Experiment settings
E.1 Common settings
Unless otherwise stated, our Gibbs-CSMC method applies the explicit backward sampling in Section 3.1 and the killing conditional resampling (Karppinen et al.,, 2023), and our PMCMC method applies the PCN proposal of Cotter et al., (2013) and stratified resampling (Chopin and Papaspiliopoulos,, 2020, Chap. 9).
For image-related experiments, we use a standard UNet commonly used in denoising diffusion models (see Ho et al.,, 2020) with three downsampling and three upsampling layers, and we use the pixel shuffle method for upsampling in the UNet. The initial number of convolution features is 64, followed by 128, and 256 in the three layers. The detailed construction of the neural network is given in our companion code repository at https://github.com/zgbkdlm/fbs.
We randomly split the MNIST dataset into 60,000 training and 10,000 test data points. As for the CelebaHQ dataset (Liu et al.,, 2015), we split it into 29,000 training and 1,000 test data points. All the image data are normalised within the range $[0,1]$.
All experiments are implemented in JAX (Bradbury et al.,, 2018) with the Flax (Heek et al.,, 2023) neural network backend. The experiments are conducted on an NVIDIA A100 80G GPU and the code required to reproduce our empirical findings can be found at https://github.com/zgbkdlm/fbs.
E.2 High-dimensional synthetic conditional sampling
@@
While the CSGM method needs the likelihood distribution $p(y\mid x_{t})$ to work (Song et al.,, 2021), in this Gaussian special case, the likelihood is analytically tractable, and hence the method can be applied exactly. For the twisted PF, we apply the GP observation model as the twisting function as per Wu et al., (2023, Eq. (8)).
@@
Figure 6 shows the trace of the PMCMC chain of one marginal variable. We see that the parameter $\delta$ affects the autocorrelation of the MCMC chain, which is also reflected in Table 4.1 and Figure 1. The chain appears to be more effective when $\delta$ is small, meaning that the measurement path $t\mapsto Y_{t}$ in Section 3.2 ideally should not move too much. Furthermore, we also see that the chain seems to gradually degenerate as the chain iteration goes for larger moves in the $Y$ space. This is a well-known defect of pseudo-marginal methods (Andrieu and Roberts,, 2009) in general, and comes from the fact that proposed paths are likely to be accepted originally due to the variance in the acceptance ratio, and then unlikely once the process has stabilised if the proposed path is too far from the current one. Solving this problem is an interesting avenue of future work and likely involves redesigning the proposal kernel for the path $Y$ to take into account the specific structure of the model.
In Section 3.1 we mentioned the memory cost due to storing the path if not implementing the online simulation of the bridge. In this experiment, we used 200 time steps, and the memory exploded when increased the number of steps to 20000 on a computer with 32GB RAM.
E.3 Conditional sampling in non-separable noising processes
@@
To show the effect of using a non-separable noising process, we conduct quantitative and qualitative experiments using Schrödinger bridges. In the quantitative experiment, we apply the same GP model in Appendix E.2, but we scale the observation noise covariance to $0.1$ to increase the correlation between $X$ and $Y$. We use a Wishart distribution to randomly generate the reference covariance matrix $Q=q\otimes q$, where $\otimes$ is the outer product, and $q\sim\mathcal{N}(0,I_{d})$. We then consider a Gaussian Schrödinger bridge between $\pi(x,y)$ and $\mathcal{N}(0,Q)$, with reference process $\mathop{}\!\mathrm{d}Z_{t}=\mathop{}\!\mathrm{d}W_{t}$ given as the standard Brownian motion, the formulation of which is analytically tractable (Bunne et al.,, 2023).
For the qualitative experiments on MNIST images, we train a Schrödinger bridge with reference process $\mathop{}\!\mathrm{d}Z_{t}=-0.5\,\beta_{t}\,Z_{t}\mathop{}\!\mathrm{d}t+\sqrt{\beta_{t}}\mathop{}\!\mathrm{d}W_{t}$ at $t\in[0,0.5]$, where $\beta_{t}\coloneqq(b_{\mathrm{max}}-b_{\mathrm{min}})\,/\,(T-t_{0})\,t+(b_{\mathrm{min}}\,T-b_{\mathrm{max}}\,t_{0})\,/\,(T-t_{0})$, and we set $b_{\mathrm{max}}=5$, $b_{\mathrm{max}}=0.02$, $t_{0}=0$, and $T=0.5$. The reference distribution $\pi_{\mathrm{ref}}$ at $T$ is a standard unit Gaussian. The number of Schrödinger bridge iterations is 20. At each Schrödinger bridge iteration, we run 20 epochs training the forward and backward models, with spatial batch size 64 and temporal batch size 32. We use the Adam (Kingma and Ba,, 2015) optimiser with learning rate decaying from $2\times 10^{-4}$ to $2\times 10^{-6}$ in a cosine curve (Loshchilov and Hutter,, 2017). For more details on the algorithm used to train the bridge, see De Bortoli et al., (2021, Algorithm 1).
@@
As shown in Figure 7, the initialisation of $X_{0}$ indeed affects the quality of PF samples due to the non-separability of the Schrödinger bridge, while our Gibbs-CSMC sampler returns valid samples. This visible bias stems from the fact that, under non-separability, the choice of $X_{0}$ has a direct impact on the distribution of the path $Y_{1:K}\mid Y_{0}=y,X_{0}$, as conditional independence is then not verified anymore. For the path $Y_{1:K}\mid\{Y_{0}=y,X_{0}\}$ to be marginally distributed according to the forward noising diffusion (as required by Trippe et al.,, 2023), one would need the initial value of $X_{0}$ to be distributed according to $\pi(x\mid y)$, which is the problem we are trying to solve and is therefore not feasible in practice. While it is reasonable to assume that approximate samples from $\pi(x\mid y)$ can be obtained via other methods, this would incur additional computation and still not eliminate the bias. On the other hand, our Gibbs-CSMC method provides a natural way to produce samples that iteratively converge to $\pi(x\mid y)$, explaining its improved performance.
@@
E.4 Image inpainting and super-resolution
@@
We first pre-train the unconditional model to approximate the resulting score function using the standard denoising score matching method of Song et al., (2020). The optimiser we used for training is Adam with a decaying cosine learning rate schedule, i.e., the learning rate starts from $2\times 10^{-4}$ and ends at $2\times 10^{-6}$ following a cosine curve (Loshchilov and Hutter,, 2017). The spatial and temporal batch sizes are 256. We also apply gradient clipping and exponential moving average (with 0.99 decay rate per two iterations) as suggested in Song and Ermon, (2020). We run the training for 3,000 epochs and use the model obtained in the last epoch.
In the inpainting task, we apply rectangles of sizes 15 and 32 for MNIST and CelebAHQ, respectively, spawning at random locations (uniformly) in the image. As for super-resolution, we use rates 4 and 2 for MNIST and CelebAHQ. Precisely, rate 2 means that for an image of size 32 by 32, we partition it into 16 by 16 squares, where each square has 2 by 2 pixels; at each square, we randomly select one pixel for the low-resolution image. For each test (inpainting/super-resolution) image, we generate 100 restored image samples, and report the statistics as in Tables 3 and 3. In the table, the LPIPS score uses an Alex net (Zhang et al.,, 2018). The commonly used Fréchet inception distance is not applied here, since it is not clear how to use it to benchmark MCMC samples (cf. the arguments in Cardoso et al.,, 2024).
For the CSGM method, we implement the conditional drift as per Song et al., (2021, Sec. I.2). As for the TPF, we also implement the same twisting function as in Wu et al., (2023, Sec. 3.3).
We show samples of inpainting and super-resolution tasks in Figures 912, to qualitatively compare the methods. We find from these figures that our Gibbs-CSMC and PMCMC methods indeed outperform all other methods, in terms of quality and correctness. [^4] While there are qualitative scoring methods (e.g., Wang et al.,, 2023; Ke et al.,, 2021), we here focus on reference-based quantitative comparisons. By correctness, we mean that the samples generated by our methods are realistic and closer to the true images. As an example, in the first panel of Figure 9, the MCMC chains generate a variety of digits “4” which indeed correspond to the true digit, while other methods can generate wrong, albeit more diverse digits.
@@
Appendix F Particle filtering coalescence
Lastly, we would also like to discuss a coalescence effect of the particle filter approach for conditional sampling (Trippe et al.,, 2023; Dou and Song,, 2024). This effect is reflected in Figure 13, where we see that all particles collapse to a single point at the terminal time of the filtering pass. Consequently, this means that it is hard to treat all the particles at the terminal time as “independent” samples (after resampling). More precisely, to draw $J$ independent samples, we would have to run the PF $J$ times independently, instead of running the PF once and using all the terminal particles as approximately the independent samples.
Figure 13: The coalescence effect of particle filtering (with 10 particles) on MNIST inpainting. The left figure shows the trajectories of the 10 particles in the filtering pass, while the right figure shows the two standard deviation of all particles for each dimension. We see that all the particles collapse to one single point at the terminal time, and this holds true across all the state dimensions.